1. The expected value of the returns for asset A is ________. Enter your answer as a percent rounded to 1 decimal, do not include the percent sign.

Asset A
Returns (%) 10 Possible Outcomes Pessimistic Most likely Optimistic Probability 0.25 0.45 0.30

2. Given the returns of two stocks J and K in the table below over the last 4 years. Find the expected return of a portfolio of 40% of stock J and 60% in stock K. Enter your answer a percent, round to one decimal, do not include the % sign.

Stock J Stock K
2015 10% 9%
2016 12% 8%
2017 13% 10%
2018 15% 11%
Answer #1

1.

Expected Return = ΣPiRi, where Pi is the probability and Ri is the return for outcome i

Hence, Expected Return = 0.25*10 + 0.45*12 + 0.30*16 = 12.7%

2.

Let return for Year n be Rn

Hence, Return for Stock J = [(1+R1)(1+R2)(1+R3)(1+R4)]1/4 - 1 = [(1+0.10)(1+0.12)(1+0.13)(1+0.15)]1/4 - 1 = 0.1249

Return for Stock K = [(1+R1)(1+R2)(1+R3)(1+R4)]1/4 - 1 = [(1+0.09)(1+0.08)(1+0.10)(1+0.11)]1/4 - 1 = 0.0949

Investment in J = 40% and Investment in K = 60%

Hence, Expected Return = ΣPiRi = 0.40*0.1249 + 0.60*0.0949 = 0.1069 or 10.69%